OpenStudy (anonymous):

Find an equation of the tangent line to: y=1/Sqrt[x] at x=9 ?

OpenStudy (anonymous):

Slope, tangent, derivative. ;o

OpenStudy (richyw):

\[y=\frac{1}{\sqrt{x}}\]\[y=x^{-1/2}\]\[\frac{dy}{dx}=-\frac{1}{2}x^{-3/2}\]

OpenStudy (anonymous):

Find the derivative of the function first, and then substitute the value of x to find the slope of the tangent line. To form an equation of a line you need a slope and a point--you know the x-value fr the point, all you need is the y. ^^ Then boom, y=m(x-p)+q and you have your equation